# Abelian congruences and similarity in varieties with a weak difference term

### v1

@@ 6. Abelian minimal congruences
(1) Suppose $(a,b)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a^{\prime},b^{\prime})$. Applying compatibility of $f$ to $(a,b)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a^{\prime},b^{\prime})$ gives $(f(a),f(b))\stackrel{{\scriptstyle\Delta}}{{\equiv}}(f(a^{\prime}),f(b^{\prime}))$, so $f^{{\mathbf{D}}}$ is well-defined. Write $f(x)=t(x,c_{1},\ldots,c_{m})$ where $t(x,\mathbf{y})$ is a term. Replacing each $c_{i}$ with $(c_{i},c_{i})/\Delta$ produces a polynomial of ${\mathbf{D}}$ which is $f^{{\mathbf{D}}}$. If $x\in D^{o}$ then $f^{{\mathbf{D}}}(x)=t(x,(c_{1},c_{1})/\Delta,\ldots,(c_{m},c_{m})/\Delta)\in D^{o}$ as each $(c_{i},c_{i})/\Delta$ is in $D^{o}$ and $D^{o}$ is a subuniverse of ${\mathbf{D}}$ by Corollary 4.8(1).
(2) Note first that if $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ and $f(C_{K})\subseteq C_{J}$, then obviously $f^{{\mathbf{D}}}(\overline{0}_{K})\in E_{J}$. Since $f^{{\mathbf{D}}}(\overline{0}_{K})\in D^{o}$ by (1), we get $f^{{\mathbf{D}}}(\overline{0}_{K})\in E_{J}\cap D^{o}=\{\overline{0}_{J}\}$, proving $f^{{\mathbf{D}}}(\overline{0}_{K})=\overline{0}_{J}$. This explains one inclusion.
To prove the other, suppose $g\in\operatorname{Pol}_{1}({\mathbf{D}})$ with $g(\overline{0}_{K})=\overline{0}_{J}$. Pick a term $t(x,\mathbf{y})$ and parameters $(u_{1},v_{1})/\Delta,\ldots,(u_{m},v_{m})/\Delta$ defining $g$. Then for every $a\in C_{K}$ and $e\in C_{J}$,
$(e,e)/\Delta=\overline{0}_{J}$ $=g(\overline{0}_{K})=g((a,a)/\Delta)$
$=t((a,a)/\Delta,(u_{1},v_{1})/\Delta,\ldots,(u_{m},v_{m})/\Delta)$
$=(t(a,\mathbf{u}),t(a,\mathbf{v}))/\Delta.$
By a property of $\Delta$ (i.e., Theorem 4.6(1)), this implies $t(a,\mathbf{u})=t(a,\mathbf{v})$ for all $a\in C_{K}$. Define $f(x):=t(x,\mathbf{u})\in\operatorname{Pol}_{1}({\mathbf{A}})$. Then for all $a\in C_{K}$ and $e\in C_{J}$ we’ve shown $(f(a),f(a))/\Delta=(e,e)/\Delta$, which, by definition of $\Delta$, forces $f(a)\stackrel{{\scriptstyle\alpha}}{{\equiv}}e$ and hence $f(C_{K})\subseteq C_{J}$. Finally, for all $(a,b)/\Delta\in E_{K}$ we have $a,b\in C_{K}$, so
$g((a,b)/\Delta)$ $=t((a,b)/\Delta,(u_{1},v_{1})/\Delta,\ldots,(u_{m},v_{m})/\Delta)$
$=(t(a,\mathbf{u}),t(b,\mathbf{v}))/\Delta$
$=(t(a,\mathbf{u}),t(b,\mathbf{u}))/\Delta$ by an earlier calculation
$=(f(a),f(b)/\Delta$
$=f^{{\mathbf{D}}}((a,b)/\Delta).$
Thus $g{\upharpoonright}_{E_{K}}=f^{{\mathbf{D}}}{\upharpoonright}_{E_{K}}$ as required.
(3) Let $e=f(0_{k})\in B_{j}$ and $g(x)=d(f(x),e,0_{j})$. Clearly $g(0_{k})=0_{j}$. Given $(a,b)/\Delta\in E_{K}$, we have
$g^{{\mathbf{D}}}((a,b)/\Delta)=(g(a),g(b))/\Delta$ $=(d(f(a),e,0_{j}),d(f(b),e,0_{j}))/\Delta=f^{{\mathbf{D}}}((a,b)/\Delta),$
where the last equality is by Lemma 5.4(2). ∎
Just as we defined ${}_{\mathbb{F}}\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{i})$ for each $i\in I$, for each $J\in[I]$ we let ${}_{\overline{\mathbb{F}}}\operatorname{Grp}_{{\mathbf{D}}}(\varphi,\overline{0}_{J})$ denote the left $\overline{\mathbb{F}}$-vector space with underlying group $\operatorname{Grp}_{{\mathbf{D}}}(\varphi,\overline{0}_{J})$ and canonical left action of $\overline{\mathbb{F}}$ given by $\overline{\lambda}\cdot x=\overline{\lambda}_{J}(x)$. An important fact about this vector space is that every range of a $\theta$-class contained in $C_{J}$ is a subspace.
Lemma 6.9.
For every $J\in[I]$ and $e\in C_{J}$, $\operatorname{ran}(e/\theta)$ is a subspace of ${}_{\overline{\mathbb{F}}}\operatorname{Grp}_{{\mathbf{D}}}(\varphi,\overline{0}_{J})$.
Proof.
Let $x\in\operatorname{ran}(e/\theta)=\operatorname{ran}(\chi_{e})$, so $x=(a,e)/\Delta$ for some $a\in e/\theta$. Let $\overline{\lambda}\in\overline{\mathbb{F}}$. If $a=e$ then $x=\overline{0}_{J}$ and $\overline{\lambda}_{J}(x)=x\in\operatorname{ran}(e/\theta)$. Assume $a\neq e$. Let $\overline{\lambda}_{J}(x)=(c,u)/\Delta$. Since $\overline{\lambda}_{J}:E_{J}\rightarrow E_{J}$ we get $c,u\in C_{J}$. Pick $j\in J$ with $c,u\in B_{j}$. Using Lemma 5.7(3), obtain $k\in J$ so that $B_{i}\rightarrow B_{k}$ and $B_{j}\rightarrow B_{k}$ witnessed by $f,g\in\operatorname{Pol}_{1}({\mathbf{A}})$ respectively. This means that $f(B_{i})\subseteq B_{k}$, $g(B_{j})\subseteq B_{k}$, $(f(v),f(z))/\Delta=(v,z)/\Delta$ for all $v,z\in B_{i}$, and $(g(v),g(z))/\Delta=(v,z)/\Delta$ for all $v,z\in B_{j}$. Let $a^{\prime}=f(a)$ and $e^{\prime}=f(e)$. By Lemma 5.7(5), we may assume that $g$ was chosen satisfying $g(u)=e^{\prime}$. Under this assumption, let $c^{\prime}=g(c)$.
Thus we have $a^{\prime},c^{\prime},e^{\prime}\in B_{k}$, $(a^{\prime},e^{\prime})/\Delta=x$, and $(c^{\prime},e^{\prime})/\Delta=\overline{\lambda}_{J}(x)$. Since $a\neq e$, we get $a^{\prime}\neq e^{\prime}$ by a property of $\Delta$ (Theorem 4.6(1)). As $(a,e),(a^{\prime},e^{\prime})\in\theta$ and $\theta$ is abelian and minimal, there exists $h\in\operatorname{Pol}_{1}({\mathbf{A}})$ satisfying $h(e^{\prime})=e$ and $h(a^{\prime})=a$ by Corollary 3.2. Then $h(B_{k})\subseteq B_{i}$, so $h(C_{J})\subseteq C_{J}$. Thus $h^{{\mathbf{D}}}{\upharpoonright}_{E_{J}}\in\overline{R}_{JJ}$ by Lemma 6.8. Hence $h^{{\mathbf{D}}}\circ\overline{\lambda}_{J}=\overline{\lambda}_{J}\circ h^{{\mathbf{D}}}{\upharpoonright}_{E_{J}}$ by condition (7b) in the definition of $\overline{\mathbb{F}}$. Also observe that $h^{{\mathbf{D}}}(x)=h^{{\mathbf{D}}}((a^{\prime},e^{\prime})/\Delta)=(h(a^{\prime}),h(e^{\prime}))/\Delta=(a,e)/\Delta=x$. Thus
$\overline{\lambda}_{J}(x)$ $=\overline{\lambda}_{J}(h^{{\mathbf{D}}}(x))$
$=h^{{\mathbf{D}}}(\overline{\lambda}_{J}(x))\qquad\qquad\mbox{as $x\in E_{J}$}$
$=h^{{\mathbf{D}}}((c^{\prime},e^{\prime})/\Delta)$
$=(h(c^{\prime}),h(e^{\prime}))/\Delta=(h(c^{\prime}),e)/\Delta\in\operatorname{ran}(\chi_{e}).$
Corollary 6.10.
Fix $\overline{\lambda}=(\overline{\lambda}_{J})_{J\in[I]}\in\overline{\mathbb{F}}$. Fix $J\in[I]$ and $i\in J$. Then for all $a,e\in B_{i}$ there exists a unique $b\in B_{i}$ satisfying $\overline{\lambda}_{J}((a,e)/\Delta)=(b,e)/\Delta$.
Proof.
Existence follows from the proof of Lemma 6.9. Uniqueness follows from injectivity of $\chi_{e}$ (Lemma 5.2). ∎
This leads to a natural action of $\overline{\mathbb{F}}$ on each group $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ in ${\mathbf{A}}$.
Definition 6.11.
Given $\overline{\lambda}=(\overline{\lambda}_{J})_{J\in[I]}\in\overline{\mathbb{F}}$ and $e\in A$, choose $J\in[I]$ so that $e\in C_{J}$ and define $\overline{\lambda}_{e}:e/\theta\rightarrow e/\theta$ so that, for each $a\in e/\theta$, $\overline{\lambda}_{e}(a)$ is the unique $b\in e/\theta$ satisfying $\overline{\lambda}_{J}((a,e)/\Delta)=(b,e)/\Delta$.
In other words, $\overline{\lambda}_{e}$ is determined by the equation $\chi_{e}\circ\overline{\lambda}_{e}=\overline{\lambda}_{J}\circ\chi_{e}$. It follows that $\overline{\lambda}_{e}=\chi_{e}^{-1}\circ\overline{\lambda}_{J}{\upharpoonright}_{\operatorname{ran}(B_{i})}\circ\chi_{e}$, which proves the next lemma.
Lemma 6.12.
Let $\overline{\lambda}\in\overline{\mathbb{F}}$ and $e\in A$. Then $\overline{\lambda}_{e}\in\operatorname{End}(\operatorname{Grp}_{{\mathbf{A}}}(\theta,e))$.
Now we define an action of $\overline{\mathbb{F}}$ on $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ in the obvious way: if $\overline{\lambda}\in\overline{\mathbb{F}}$ and $a\in e/\theta$, define $\overline{\lambda}\cdot a:=\overline{\lambda}_{e}(a)$. It can be routinely checked that this gives $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ the structure of a left $\overline{\mathbb{F}}$-vector space, which we denote by ${}_{\overline{\mathbb{F}}}\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$. We get the following improvement to Lemma 5.2 and Corollary 5.8 from Section 5.
Proposition 6.13.
Let ${\mathbf{A}}$ be an algebra in a variety with a weak difference term, let $\theta$ be an abelian minimal congruence, let $\alpha,\Delta,\varphi,D^{o},{\mathbf{D}}$ be as in Notation 6.6, and let $\overline{\mathbb{F}}$ be the division ring associated to ${\mathbf{D}},\varphi,D^{o}.$ Let $C$ be an $\alpha$-class.
  1. (1)
    For each $e\in C$, the map $\chi_{e}:a\mapsto(a,e)/\Delta$ is a left $\overline{\mathbb{F}}$-vector space embedding of ${}_{\overline{\mathbb{F}}}\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ into ${}_{\overline{\mathbb{F}}}\operatorname{Grp}_{{\mathbf{D}}}(\varphi,0_{C})$.
  2. (2)
    The set $S_{C}:=\{\operatorname{ran}(B)\,:\,\text{$B\in A/\theta$, $B\subseteq C$}\}$ is a directed set of subspaces of ${}_{\overline{\mathbb{F}}}\operatorname{Grp}_{{\mathbf{D}}}(\varphi,0_{C})$ whose union is ${}_{\overline{\mathbb{F}}}\operatorname{Grp}_{{\mathbf{D}}}(\varphi,0_{C})$.
Recall that we also have the left $\mathbb{F}$-vector space ${}_{\mathbb{F}}\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ where $\mathbb{F}$ is the division ring associated to ${\mathbf{A}}$, $\theta$ and a transversal $T$ for $\theta$ which we fixed at the beginning of the section. Let $(0_{i})_{i\in I}$ be as in Notation 6.1. In the remainder of this section will show that ${}_{\mathbb{F}}\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ and ${}_{\overline{\mathbb{F}}}\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ are essentially the same.
Definition 6.14.
Define $\Psi:\overline{\mathbb{F}}\rightarrow\prod_{i\in I}\operatorname{End}(\operatorname{Grp}_{{\mathbf{A}}}(\theta,0_{i}))$ by $\Psi(\overline{\lambda})=(\overline{\lambda}_{0_{i}})_{i\in I}$.
Lemma 6.15.
Let ${\mathbf{A}},\theta,T,\alpha,\Delta,{\mathbf{D}},D^{o}$ be as in Notation 6.6. Let $\mathbb{F}$ be the division ring associated to ${\mathbf{A}},\theta,T$ and let $\overline{\mathbb{F}}$ be the division ring associated to ${\mathbf{D}},\varphi,D^{o}$.
  1. (1)
    $\operatorname{ran}(\Psi)=\mathbb{F}$.
  2. (2)
    $\Psi$ is a ring isomorphism from $\overline{\mathbb{F}}$ to $\mathbb{F}$.
  3. (3)
    Fix $e\in A$. If $\cdot$ denotes the action of $\overline{\mathbb{F}}$ on $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ and $\cdot^{e}$ denotes the action of $\mathbb{F}$ on $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$, then for all $\overline{\lambda}\in\overline{\mathbb{F}}$ and $a\in e/\theta$, $\overline{\lambda}\cdot a=\Psi(\overline{\lambda})\cdot^{e}a$.
Proof.
(1) Let $\overline{\lambda}=(\overline{\lambda}_{J})_{J\in[I]}\in\overline{\mathbb{F}}$, so $\Psi(\overline{\lambda})=(\overline{\lambda}_{0_{i}})_{i\in I}$. To show $\Psi(\overline{\lambda})\in\mathbb{F}$, we must verify condition (2b) in Notation 6.1. Let $j,k\in I$ and $r\in R_{jk}$; thus $r=f{\upharpoonright}_{B_{k}}$ for some $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ with $f(0_{k})=0_{j}$. Let $J=[j]$ and $K=[k]$. Then $f(C_{K})\subseteq C_{J}$, so $f^{{\mathbf{D}}}{\upharpoonright}_{E_{K}}\in\overline{R}_{JK}$. Condition (7b) in Notation 6.6 gives $\overline{\lambda}_{J}\circ f^{{\mathbf{D}}}{\upharpoonright}_{E_{K}}=f^{{\mathbf{D}}}{\upharpoonright}_{E_{K}}\circ\overline{\lambda}_{K}$. Also note that for all $a\in B_{k}$,
$\chi_{0_{j}}(r(a))=(f(a),0_{j})/\Delta=(f(a),f(0_{k}))/\Delta=f^{{\mathbf{D}}}((a,0_{k})/\Delta)=f^{{\mathbf{D}}}(\chi_{0_{k}}(a))$
which proves $\chi_{0_{j}}\circ r=f^{{\mathbf{D}}}{\upharpoonright}_{E_{K}}\circ\chi_{0_{k}}$. Combining this with $\chi_{0_{j}}\circ\overline{\lambda}_{0_{j}}=\overline{\lambda}_{J}\circ\chi_{0_{j}}$ gives
$\chi_{0_{j}}\circ\overline{\lambda}_{0_{j}}\circ r=\overline{\lambda}_{J}\circ\chi_{0_{j}}\circ r=\overline{\lambda}_{J}\circ f^{{\mathbf{D}}}{\upharpoonright}_{E_{K}}\circ\chi_{0_{k}}$
and similarly
$\chi_{0_{j}}\circ r\circ\overline{\lambda}_{0_{k}}=f^{{\mathbf{D}}}{\upharpoonright}_{E_{K}}\circ\chi_{0_{k}}\circ\overline{\lambda}_{0_{k}}=f^{{\mathbf{D}}}{\upharpoonright}_{E_{K}}\circ\overline{\lambda}_{K}\circ\chi_{0_{k}}.$
Since $\overline{\lambda}_{J}\circ f^{{\mathbf{D}}}{\upharpoonright}_{E_{K}}=f^{{\mathbf{D}}}{\upharpoonright}_{E_{K}}\circ\overline{\lambda}_{K}$ and $\chi_{0_{j}}$ is injective (Lemma 5.2), we get $\overline{\lambda}_{0_{j}}\circ r=r\circ\overline{\lambda}_{0_{k}}$ as required by condition (2b) in Notation 6.1. This proves $\Psi(\overline{\lambda})\in\mathbb{F}$, so $\operatorname{ran}(\Psi)\subseteq\mathbb{F}$.
Towards proving the opposite inclusion, we first show the following.
Claim 6.15.1.
Suppose $\lambda=(\lambda_{i})_{i\in I}\in\mathbb{F}$. Then for all $J\in[I]$ and $i,j\in J$, and all $a\in B_{i}$ and $a^{\prime}\in B_{j}$, if $(a,0_{i})\stackrel{{\scriptstyle\Delta}}{{\equiv}}(a^{\prime},0_{j})$, then $(\lambda_{i}(a),0_{i})\stackrel{{\scriptstyle\Delta}}{{\equiv}}(\lambda_{j}(a^{\prime}),0_{j})$.
Proof of Claim 6.15.1.
Using Lemma 5.7(3), choose $k\in J$ with $B_{i},B_{j}\rightarrow B_{k}$ witnessed by $f,g\in\operatorname{Pol}_{1}({\mathbf{A}})$ respectively. By Lemma 5.7(5), we can assume that $f(0_{i})=g(0_{j})=0_{k}$. Thus $(f(a),0_{k})/\Delta=(a,0_{i})/\Delta=(a^{\prime},0_{j})/\Delta=(g(a^{\prime}),0_{k})/\Delta$ by the definition of $\rightarrow$, and so $f(a)=g(a^{\prime})$ by injectivity of $\chi_{0_{k}}$. Let $r=f{\upharpoonright}_{B_{i}}$ and $s=g{\upharpoonright}_{B_{j}}$. Then $r\in R_{ki}$ and $s\in R_{kj}$. Then
$(\lambda_{i}(a),0_{i})/\Delta$ $=(r(\lambda_{i}(a)),0_{k})/\Delta$ as $r=f{\upharpoonright}_{B_{i}}$ and $f$ witnesses $B_{i}\rightarrow B_{k}$
$=(\lambda_{k}(r(a)),0_{k})/\Delta$ as $\lambda_{k}\circ r=r\circ\lambda_{i}$
and similarly $(\lambda_{j}(a^{\prime}),0_{j})/\Delta=(\lambda_{k}(s(a^{\prime})),0_{k})$. As $r(a)=s(a^{\prime})$, we get $(\lambda_{i}(a),0_{i})/\Delta=(\lambda_{j}(a^{\prime}),0_{j})/\Delta$ as required. ∎
Observe that for each $J\in[I]$, every element of $E_{J}$ can be written in the form $(a,0_{j})/\Delta$ for some $j\in J$ and $a\in B_{j}$. This observation, combined with Claim 6.15.1, justifies the following definition.
Definition 6.15.2.
For $\lambda=(\lambda_{i})_{i\in I}\in\mathbb{F}$ and $J\in[I]$, define $\lambda_{J}:E_{J}\rightarrow E_{J}$ as follows: if $i\in J$ and $a\in B_{i}$, then $\lambda_{J}((a,0_{i})/\Delta):=(\lambda_{i}(a),0_{i})/\Delta$.
In other words, $\lambda_{J}$ is determined by the equations $\lambda_{J}\circ\chi_{0_{i}}=\chi_{0_{i}}\circ\lambda_{i}$ for all $i\in J$. It follows that for each $i\in J$, $\lambda_{J}{\upharpoonright}_{\operatorname{ran}(B_{i})}=\chi_{0_{i}}\circ\lambda_{i}\circ\chi_{0_{i}}^{-1}$. Since $\operatorname{Grp}_{{\mathbf{D}}}(\varphi,\overline{0}_{J})$ is the union of the directed family $\{\operatorname{ran}(B_{i})\,:\,\text{$i\in J$}\}$ of subgroups by Corollary 5.8, it follows that $\lambda_{J}\in\operatorname{End}(\operatorname{Grp}_{{\mathbf{D}}}(\varphi,\overline{0}_{J}))$.
Definition 6.15.3.
Define $\Phi:\mathbb{F}\rightarrow\prod_{J\in[I]}\operatorname{End}(\operatorname{Grp}_{{\mathbf{D}}}(\varphi,\overline{0}_{J}))$ as follows: given $\lambda\in\mathbb{F}$, set $\Phi(\lambda):=(\lambda_{J})_{J\in[I]}$.
Claim 6.15.4.
$\operatorname{ran}(\Phi)\subseteq\overline{\mathbb{F}}$ and $\Psi\circ\Phi=\mathrm{id}_{\mathbb{F}}$. Thus $\operatorname{ran}(\Psi)=\mathbb{F}$.
Proof of Claim 6.15.4.
Let $\lambda=(\lambda_{i})_{i\in I}\in\mathbb{F}$. To prove $\Phi(\lambda)\in\overline{\mathbb{F}}$, we must show $\lambda_{J}\circ\overline{r}=\overline{r}\circ\lambda_{K}$ for all $J,K\in[I]$ and all $\overline{r}\in\overline{R}_{JK}$. Fix $J,K\in[I]$ and $\overline{r}\in\overline{R}_{JK}$. By Corollary 5.8, it is enough to prove $\lambda_{J}\circ\overline{r}{\upharpoonright}_{\operatorname{ran}(B_{k})}=\overline{r}\circ\lambda_{K}{\upharpoonright}_{\operatorname{ran}(B_{k})}$ for all $k\in K$. Fix $k\in K$. Choose $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ with $f(C_{K})\subseteq C_{J}$ and such that $\overline{r}=f^{{\mathbf{D}}}{\upharpoonright}_{E_{K}}$. By Lemma 6.8(3), we can assume $f(0_{k})=0_{j}$ for some $j\in J$. Define $r:=f{\upharpoonright}_{B_{k}}$. Then $r\in R_{jk}$. Now for any $a\in B_{k}$,
$\lambda_{J}(\overline{r}((a,0_{k})/\Delta))$ $=\lambda_{J}((r(a),0_{j})/\Delta)$
$=(\lambda_{j}(r(a)),0_{j})/\Delta$ by definition of $\lambda_{J}$
$=(r(\lambda_{k}(a)),0_{j})/\Delta$ as $\lambda_{j}\circ r=r\circ\lambda_{k}$
$=\overline{r}(\lambda_{K}((a,0_{k})/\Delta))$ by definition of $\mbox{by definition of $\lambda_{K}$}.$
This proves $\Phi(\lambda)\in\overline{\mathbb{F}}$ and hence $\operatorname{ran}(\Phi)\subseteq\overline{\mathbb{F}}$. Again let $\lambda=(\lambda_{i})_{i\in I}\in\mathbb{F}$ and let $\overline{\lambda}=(\lambda_{J})_{J\in[I]}=\Phi(\lambda)\in\overline{\mathbb{F}}$. To show $\Psi(\overline{\lambda})=\lambda$, we must show $\overline{\lambda}_{0_{i}}=\lambda_{i}$ for each $i\in I$. Fix $J\in[I]$ and $i\in J$. We have $\lambda_{J}\circ\chi_{0_{i}}=\chi_{0_{i}}\circ\lambda_{i}$ by the definition of $\lambda_{J}$. But this means that $\lambda_{i}$ satisfies the equation which determines $\overline{\lambda}_{0_{i}}$ (see the discussion following Definition 6.11). This proves $\Psi(\overline{\lambda})=\lambda$ and hence $\Psi\circ\Phi=\mathrm{id}_{\mathbb{F}}$, which completes the proof of Claim 6.15.4 and of (1). ∎
(2) It is a tedious but mostly routine exercise to show that $\Psi$ is a unital ring homomorphism. For example, to prove that $\Psi$ preserves addition, it suffices to prove that if $\overline{\lambda},\overline{\mu},\overline{\tau}\in\overline{\mathbb{F}}$ with $\overline{\lambda}=(\overline{\lambda}_{J})_{J\in[I]}$ etc. and such that $\overline{\tau}=\overline{\lambda}+\overline{\mu}$ calculated in $\overline{\mathbb{F}}$, then for all $e\in A$ and $a\in e/\theta$ we have $\overline{\tau}_{e}(a)=\overline{\lambda}_{e}(a)+\overline{\mu}_{e}(a)$ calculated in $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$. The following calculation proves this. Given $J\in[I]$, $e\in C_{J}$ and $a\in e/\theta$,
$\chi_{e}(\overline{\tau}_{e}(a))$ $=\overline{\tau}_{J}(\chi_{e}(a))$ $\chi_{e}\circ\overline{\tau}_{e}=\overline{\tau}_{J}\circ\chi_{e}$, by definition of $\overline{\tau}_{e}$
$=\overline{\lambda}_{J}(\chi_{e}(a))+\overline{\mu}_{J}(\chi_{e}(a))$ definition of $\overline{\lambda}+\overline{\mu}$ in $\overline{\mathbb{F}}$
$=\chi_{e}(\overline{\lambda}_{e}(a))+\chi_{e}(\overline{\mu}_{e}(a))$ definitions of $\overline{\lambda}_{e}$ and $\overline{\mu}_{e}$
$=\chi_{e}(\overline{\lambda}_{e}(a)+\overline{\mu}_{e}(a))$ as $\chi_{e}$ is a group homomorphism.
Since $\chi_{e}$ is injective, we get $\overline{\tau}_{e}(a)=\overline{\lambda}_{e}(a)+\overline{\tau}_{e}(a)$ as required. The proofs that $\Psi$ preserves multiplication and the unit are no harder. Thus $\Psi$ is a surjective unital ring homomorphism, and since $\overline{\mathbb{F}}$ is a division ring and $\mathbb{F}$ is nontrivial, $\Psi$ must be an isomorphism.
(3) Let $\overline{\lambda}=(\overline{\lambda}_{J})_{J\in[I]}\in\overline{\mathbb{F}}$, $e\in A$, and $a\in e/\theta$. Let $i\in I$ so that $a,e\in B_{i}$ and let $J=[i]$. Let $\cdot$ denote the action of $\overline{\mathbb{F}}$ on $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ and let $\cdot^{e}$ denote the action of $\mathbb{F}$ on $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$. Let $b=d(a,e,0_{i})$, $x=\overline{\lambda}_{0_{i}}(b)$, and $c=d(x,0_{i},e)$. Thus by definition,
$(x,0_{i})/\Delta=\overline{\lambda}_{J}((b,0_{i})/\Delta)\quad\mbox{and}\quad\Psi(\overline{\lambda})\cdot^{e}a=c.$
On the other hand, Lemma 5.4(1) gives
$(a,e)/\Delta=(b,0_{i})/\Delta\quad\mbox{and}\quad(x,0_{i})/\Delta=(c,e)/\Delta.$
Thus $(c,e)/\Delta=\overline{\lambda}_{J}((a,e)/\Delta)$, which implies $\overline{\lambda}\cdot a=\overline{\lambda}_{e}(a)=c$ as required. ∎
In summary, we have proved the following.
Proposition 6.16.
[…]

### v2

@@ 6.2. The general case
Observe that $(0:\theta/\gamma)=\alpha/\gamma$ by Lemma 2.3(1), so we may assume with no loss of generality that $\gamma=0$. Let $\overline{{\mathbf{A}}}$, $\overline{\varepsilon}$ and $\overline{\alpha}$ denote ${\mathbf{A}}/\delta$, $\varepsilon/\delta$ and $\alpha/\delta$ respectively. Note that $(0:\overline{\varepsilon})=\overline{\alpha}$ by Lemma 2.3(1) and Proposition 3.14, so $D(\overline{{\mathbf{A}}},\overline{\varepsilon})=\overline{{\mathbf{A}}}(\overline{\varepsilon})/\Delta_{\overline{\varepsilon},\overline{\alpha}}$.
(1) Let $\psi:{\mathbf{A}}(\theta)\rightarrow\overline{{\mathbf{A}}}(\overline{\varepsilon})$ be the homomorphism given by by $\psi((a,b))=(a/\delta,b/\delta)$, let $\nu:\overline{{\mathbf{A}}}(\overline{\varepsilon})\rightarrow D(\overline{{\mathbf{A}}},\overline{\varepsilon})$ be the natural projection map, and let $\Psi=\nu\circ\psi$. Since $\varepsilon=\delta\circ\theta\circ\delta$ by Proposition 3.14, $\psi$ is surjective, so $\Psi$ is also surjective. We must show that $\ker(\Psi)=\Delta_{\theta,\alpha}$. Let $\Delta=\Delta_{\theta,\alpha}$. Because $\varepsilon=\delta\circ\theta\circ\delta$, we have
(6.3) $M(\overline{\varepsilon},\overline{\alpha})=\left\{\begin{pmatrix}a/\delta&a^{\prime}/\delta\\ b/\delta&b^{\prime}/\delta\end{pmatrix}:\begin{pmatrix}a&a^{\prime}\\ b&b^{\prime}\end{pmatrix}\in M(\theta,\alpha)\right\}.$
Hence $\Delta\subseteq\ker(\Psi)$. Before proving the opposite inclusion, we establish the following.
Claim. If $(a,b),(r,s)\in\theta$ and $(a,r),(b,s)\in\delta$, then $(a,b)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(r,s)$.
Proof of Claim.
Let $d$ be a weak difference term for $\mathcal{V}$. Note first that $d(a,r,r)\stackrel{{\scriptstyle\theta}}{{\equiv}}d(b,s,r)$ and $d(a,r,r)\stackrel{{\scriptstyle\delta}}{{\equiv}}d(a,a,a)=a=d(b,b,a)\stackrel{{\scriptstyle\delta}}{{\equiv}}d(b,s,r)$. Since $\theta\cap\delta=0$, this proves
(6.4) $d(a,r,r)=d(b,s,r).$
Next, note that
(6.5) $d(\underline{r},r,r)=r=d(\underline{r},s,s),$
so since $(r,s)\in\theta$ and $(r,b)\in\delta\vee\theta=\varepsilon\leq\alpha$ and $C(\alpha,\theta;0)$, we can replace the underlined occurrences of “rr” in equation 6.5 by “bb” to get
(6.6) $d(b,r,r)=d(b,s,s).$
Now we build some $(\theta,\alpha)$-matrices. Let
$M_{1}=\begin{pmatrix}a&a\\ b&b\end{pmatrix},~~M_{2}=\begin{pmatrix}a&r\\ a&r\end{pmatrix},~~M_{3}=\begin{pmatrix}b&s\\ b&s\end{pmatrix},~~M_{4}=\begin{pmatrix}s&s\\ s&s\end{pmatrix},~~M_{5}=\begin{pmatrix}r&r\\ s&s\end{pmatrix}$.
Clearly $M_{1},\ldots,M_{5}\in M(\theta,\alpha)$. Applying the weak difference term gives
$N:=d(M_{1},M_{2},M_{2})$ $=\begin{pmatrix}a&d(a,r,r)\\ b&d(b,r,r)\end{pmatrix}\in M(\theta,\alpha)$
$N^{\prime}:=d(M_{3},M_{4},M_{5})$ $=\begin{pmatrix}d(b,s,r)&r\\ d(b,s,s)&s\end{pmatrix}\in M(\theta,\alpha).$
The second column of $N$ and the first column of $N^{\prime}$ are identical by equations (6.4) and (6.6). Since $\Delta$ is the “horizontal transitive closure” of $M(\theta,\alpha)$, this proves $(a,b)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(r,s)$, proving the Claim. ∎
Now we prove $\ker(\Psi)\subseteq\Delta$. Assume $((a,b),(a^{\prime},b^{\prime}))\in\ker(\Psi)$; that is,
$(a,b),(a^{\prime},b^{\prime})\in\theta$ and $(a/\delta,b/\delta)\stackrel{{\scriptstyle\Delta_{\overline{\varepsilon},\overline{\alpha}}}}{{\equiv}}(a^{\prime}/\delta,b^{\prime}/\delta)$.
By definition of $\Delta_{\overline{\varepsilon},\overline{\alpha}}$ and equation 6.3, there exist $n\geq 0$ and
(6.7) $\begin{pmatrix}r_{i}&a_{i}\\ s_{i}&b_{i}\end{pmatrix}\in M(\theta,\alpha)\quad\mbox{for $i\leq n$}$
[…]